A random sample of ten measurements were obtained from a normally distributed population with mean u=6.5. The sample values are X-4.2 and s 2.




a. Test the null hypothesis that the mean of the population against the alternative hypothesis, μ = 6.5. Use a = 0.05.




b. Test the null hypothesis that the mean of the population against the alternative hypothesis, u 6.5. Use a = 0.05

1
Expert's answer
2022-05-16T15:55:51-0400

a.

The following null and alternative hypotheses need to be tested:

"H_0:\mu=6.5"

"H_1:\mu>6.5"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a right-tailed test is "z_c = 1.6449."

The rejection region for this right-tailed test is "R = \{z:z> 1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{4.2-6.5}{2/\sqrt{10}}\approx-3.6366"

Since it is observed that "z=-3.6366<1.6449=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=P(z>-3.6366)=0.999862," and since "p=0.999862>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is greater than 90, at the "\alpha = 0.05" significance level.


b.

The following null and alternative hypotheses need to be tested:

"H_0:\mu=6.5"

"H_1:\mu\not=6.5"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{4.2-6.5}{2/\sqrt{10}}\approx-3.6366"

Since it is observed that "z =- 3.6366 <1.96=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(z<-3.6366)=0.000276," and since "p=0.000276<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 6.5, at the "\alpha = 0.05" significance level.